OpenStudy (anonymous):

Given the function f(x) = log3(x + 4), find the value of f−1(3). f−1(3) = 7 f−1(3) = 12 f−1(3) = 23 f−1(3) = 31

OpenStudy (aum):

\(\Large f(x) = \log_3(x+4)\) ?

OpenStudy (anonymous):

yes

OpenStudy (aum):

In y = f(x), to evaluate, y when x = 3, you would put x = 3 in f(x) and calculate y. But to to calculate f^-1(3), you set y = 3 and solve for x. Set x = each answer choice in f(x) and see which one gives you 3.

OpenStudy (aum):

\(\Large f(x) = \log_3(x+4) \) \(\Large f^{-1}(3) ~~\text{ means } ~~~3 = \log_3(x+4)\) Try each answer choice for x and see which one gives 3.

OpenStudy (aum):

\( \Large 3 = \log_3(x+4) \\ \Large 3^3 = x + 4 \\ \Large 27 = x + 4 \\ \Large x = 23 \\ \Large f^{-1}(3) = 23 \)

OpenStudy (anonymous):

Thankyou!!